JACKIE WANTS TO MAKE A MIXTURE OF NUTS TO SELL IN HER STORE CONSISTING OF HAZELNUTS AND CASHEWS.HAZELNUTS COST $6.50 PER POUND AND CASHEWS COST $4.50 PER POUND. IF JACKIE WANTS 60 POUNDS TOTAL OF MIXTURE AND THE COST TO BE $5.10 PER POUND, HOW MANY POUND OF EACH WILL SHE NEED?
Answer by Earlsdon(6294)

Let x = the number of lbs. of hazelnuts at $6.50 per pound and (60-x) = the number of lbs. of cashews at $4.50 per pound.
The mixture (the sum of these two) is to be 60 lbs at $5.10 per pound.
This can be expressed algebraically by:
6.5x+4.5(60-x) = 5.1(60) Simplify.
6.5x+270-4.5x = 306 Combine like-terms.
2x+270 = 306 Subtract 270 from both sides.
2x = 36 Divide both sides by 2.
x = 18
Jackie will need to mix 18 lbs. of hazelnuts @ $6.50 per pound with (60-18 = 42)lbs. of cashews @ $4.50 per pound to obtain 60 pounds of the mixture @ $5.10 per pound.